The value of h is?
d=h+q/2

collect like terms
24/9 m + 7/9 n - 14/9 m + 12/9 m

Answer by stanbon(75887)

The value of h is?
d = (h+q)/2
2d = h+q
h = 2d-q
-----------------------
collect like terms
24/9 m + 7/9 n - 14/9 m + 12/9 m
----------
(24/9+12/9-14/9)m + (7/9)n
---
= (22/9)m + (7/9)n
------------
= (22m + 7n)/9
====================
Cheers,
Stan H.
===========